Grade 11Mechanics

Find the time of flight of a projectile thrown horizontally with speed 10m/sec from a long inclined plane which makes an angle of 45 degree from the horizontal.

Purva
8 Years agoGrade 11

2 Answers

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
We know as
Time of flight = 2u sin theta/g
 
 = 2 * 10 * sin45°/10 = 2 * 1/sqrt(2)
 
 = sqrt(2) = 1.414 sec
Akshat Gupta
ApprovedApproved Tutor Answer8 Years ago
Vertical Component of velocity =usin∂  m/s
Acceleration due to gravity=g=10m/s^2
Time of Flight = 2usin∂/g
                          = 2×10×sin45°/10
                          = 20×1/√2×1/10
                          = 2/√2=√2=1.414s